The length of a rectangle is 3 feet less than 3 times the width. If the perimeter is 146 feet, what is the length and the width of the rectangle?
Answer by checkley75(3666)

LENGTH=3W-3
WIDTH=W
PERIMETER=2*L+2*W
146=2(3W-3)+2W
146=6W-6+2W
146+6=8W
152=8W
W=152/8
W=19 IS THE WIDTH.
L=3*19-3
L=57-3
L=54 IS THE LENGTH.
PROOF
146=2(3*19-3)+2*19)
146=2(57-3)+38
146=2*54+38
146=108+38
146=146